OpenStudy (anonymous):

I want y' of this function:

OpenStudy (anonymous):

\[-\sin(x ^{2}+2y).(2x+2y \prime)=x.e ^{y ^{2}}.2y.y \prime+e ^{y ^{2}}\]

geerky42 (geerky42):

Isolating y' should be easy, apply distributive property in LHS

geerky42 (geerky42):

After this, get all term with y' to one side and all term without y' to other, then you can factor y' out

geerky42 (geerky42):

Can you do this? @soso707

OpenStudy (anonymous):

\[2y \prime.y \prime=x.e ^{y ^{2}}.2y+e ^{y ^{2}}-2x+\sin(x ^{2}+2y)\] it's true?

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

help me :(

mathslover (mathslover):

\(-\sin(x ^{2}+2y).(2x+2y \prime)=x.e ^{y ^{2}}.2y.y \prime+e ^{y ^{2}}\ \\ -2x \sin (x^{2} + 2y) -2y' (\sin (x^{2} + 2y) ) = x.e ^{y ^{2}}.2y.y \prime+e ^{y ^{2}} \) \(2y' (\sin (x^2 + 2y)) + x.e^{y^2}. 2y.y' = -2x \sin (x^2 +2y) - e^{y ^2} \) \(y' ( 2\sin (x^2 +2y) + x .e^{y^2} . 2y ) = -2x \sin (x^2 +2y) -e^{y^2} \)

mathslover (mathslover):

\(y' = \cfrac{-2x \sin (x^2 +2y) - e^{y^2} }{2\sin (x^2 +2y) + x.e^{y^2}. 2y} \) This is not so easy.. I think. Wait, I will try something new.

OpenStudy (anonymous):

\[-\sin(x ^{2}+2y).(2x+2y \prime)=x.e ^{y ^{2}}.2y.y \prime+e ^{y ^{2}}\ \\ -2x \sin (x^{2} + 2y) -2y' (\sin (x^{2} + 2y) ) = x.e ^{y ^{2}}.2y.y \prime+e ^{y ^{2}} \] how can you get this part?

mathslover (mathslover):

Using distributive property : a(b+c) = ab + ac \((2x + 2y')((-\sin (x^2 +2y)) = 2x \times \left( - \sin (x^2 + 2y) \right) + 2y' \times \left( - \sin (x^2 +2y) \right) \)

OpenStudy (anonymous):

aha

OpenStudy (anonymous):

so the fainl solution \[y' = \cfrac{-2x \sin (x^2 +2y) - e^{y^2} }{2\sin (x^2 +2y) + x.e^{y^2}. 2y} \]

mathslover (mathslover):

So, have you got it?

OpenStudy (anonymous):

yes from you

OpenStudy (anonymous):

thank you

mathslover (mathslover):

You're Welcome Soso :)